OpenStudy (adrianna.gongora):

Graph one of your 2nd degree functions from question 1. Identify which function you used and the key features of your graph. Explain how to find them algebraically.

OpenStudy (adrianna.gongora):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it. The function f(x) is a difference of squares. say a -b where a and b are squares (sqrta - sqrtb)(sqrta + sqtb) The function g(x) is a sum of squares. does not factor The function h(x) is a perfect square trinomial. this has form a^2 + 2ab + b^2 factors to (a+b)^2

OpenStudy (adrianna.gongora):

That is #1 .

OpenStudy (adrianna.gongora):

It's okay c:

OpenStudy (adrianna.gongora):

@Hayhayz Have you done this by any chance in your math class? Module 3?

Hayhayz (hayhayz):

Im bad at math lol @TheSmartOne

OpenStudy (adrianna.gongora):

It's always the one I tag you in and need help with lol :3

Hayhayz (hayhayz):

@Mehek14 @Luigi0210 @ganeshie8 @KendrickLamar2014 @jabez177 @perl Can one of you help her?

OpenStudy (kendricklamar2014):

Example: For #1 \[f(x) = x^2-9 \] \[f(x) = (x+3)(x-3) \] \[f(x) = x^2 + 9 <-- this ~ cannot ~ be ~ factored \] \[f(x) = x^2 + 6x + 9 \] \[f(x) = (x+3)^2 \] \[f(x) 2x^2 + 2x + 2 \] \[f(x) = 2(x^2 + x + 1) \]

OpenStudy (adrianna.gongora):

well this is the question I posted number one because he had to do with number one .. Graph one of your 2nd degree functions from question 1. Identify which function you used and the key features of your graph. Explain how to find them algebraically.

OpenStudy (kendricklamar2014):

I dont understand...

OpenStudy (adrianna.gongora):

the question i need help with is Graph one of your 2nd degree functions from question 1. Identify which function you used and the key features of your graph. Explain how to find them algebraically.

OpenStudy (kendricklamar2014):

Well lets go step by step. 1. Which one do you want to graph? 2. What function did you use? 3. What are the key features in your graph (once graphed) 4. How did you find them algebraically?

OpenStudy (adrianna.gongora):

1. function h(x) The function h(x) is a perfect square trinomial. this has form a^2 + 2ab + b^2 factors to (a+b)^2 is this what you wanted to know?

OpenStudy (kendricklamar2014):

Yes, that is.

OpenStudy (adrianna.gongora):

So what do I do next I don't know how to answer the question lol am I supposed to graph something?

OpenStudy (kendricklamar2014):

Yes, you need to graph one of your 2nd degree functions. Which one do you want to graph?

OpenStudy (adrianna.gongora):

2nd degree functions? could I graph any of the ones from #1 or the h(x) one?

OpenStudy (kendricklamar2014):

See, that's what im confused. Can you show me what was question 1 and your answer for that question?

OpenStudy (adrianna.gongora):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it. The function f(x) is a difference of squares. say a -b where a and b are squares (sqrta - sqrtb)(sqrta + sqtb) The function g(x) is a sum of squares. does not factor The function h(x) is a perfect square trinomial. this has form a^2 + 2ab + b^2 factors to (a+b)^2 The function j(x) can only have a GCF factored out of it. this has form ax^2 + abx + ac factors to a(x^2 + bx + c) That is the question and my answer is it wrong?

OpenStudy (kendricklamar2014):

Ahh okay. So im guessing your answer is: The function f(x) is a difference of squares. say a -b where a and b are squares (sqrta - sqrtb)(sqrta + sqtb) The function g(x) is a sum of squares. does not factor The function h(x) is a perfect square trinomial. this has form a^2 + 2ab + b^2 factors to (a+b)^2 The function j(x) can only have a GCF factored out of it. this has form ax^2 + abx + ac factors to a(x^2 + bx + c)

OpenStudy (kendricklamar2014):

Am I right?

OpenStudy (adrianna.gongora):

yes is it wrong?

OpenStudy (i<3u):

Does this help? https://answers.yahoo.com/question/index?qid=20131202153526AAtnEF7# @Adrianna.Gongora

OpenStudy (adrianna.gongora):

I already looked at that c: someone sent me that link for #1 I need help on #4 :) which is Graph one of your 2nd degree functions from question 1. Identify which function you used and the key features of your graph. Explain how to find them algebraically.

OpenStudy (i<3u):

Oh okay, let me see what I can find.

OpenStudy (adrianna.gongora):

c: Thank you

OpenStudy (kendricklamar2014):

@Adrianna.Gongora I am going to have to go offline. I need to go pickup my cousin from school. Sorry to leave but I hope @I<3U can help you.

OpenStudy (adrianna.gongora):

It's okay c: no problem thank you @KendrickLamar2014 c:

OpenStudy (kendricklamar2014):

:)

OpenStudy (i<3u):

Alright, so what are we going to graph @Adrianna.Gongora

OpenStudy (adrianna.gongora):

That's what I am trying to figure out ;-;

OpenStudy (i<3u):

Well what are your functions that you answered in question #1?

OpenStudy (adrianna.gongora):

The function f(x) is a difference of squares. say a -b where a and b are squares (sqrta - sqrtb)(sqrta + sqtb) The function g(x) is a sum of squares. does not factor The function h(x) is a perfect square trinomial. this has form a^2 + 2ab + b^2 factors to (a+b)^2 The function j(x) can only have a GCF factored out of it. this has form ax^2 + abx + ac factors to a(x^2 + bx + c)

OpenStudy (i<3u):

Alright, so which one do you want to graph?

OpenStudy (adrianna.gongora):

Which one would be easier?

OpenStudy (i<3u):

Hmm the last two are pretty simple to graph. Lets graph: a(x^2 + bx + c)

OpenStudy (i<3u):

http://prntscr.com/auswlo

OpenStudy (adrianna.gongora):

is that the graph for it?

OpenStudy (i<3u):

Yes, that is

OpenStudy (adrianna.gongora):

is that all I needed for #4?

OpenStudy (i<3u):

No, you also need to: Identify which function you used and the key features of your graph. Explain how to find them algebraically. @KendrickLamar2014 explained to you the steps to solve problem #4

OpenStudy (adrianna.gongora):

Example: For #1 f(x)=x2−9 f(x)=(x+3)(x−3) f(x)=x2+9<−−this cannot be factored f(x)=x2+6x+9 f(x)=(x+3)2 f(x)2x2+2x+2 f(x)=2(x2+x+1) this ?

OpenStudy (i<3u):

No, this: Well lets go step by step. 1. Which one do you want to graph? 2. What function did you use? 3. What are the key features in your graph (once graphed) 4. How did you find them algebraically?

OpenStudy (adrianna.gongora):

Okay so I would put 1. a(x^2 + bx + c) 2. Polynomial Function 3. what do they mean by this one for number 3? 4.

OpenStudy (i<3u):

Im trying to solve #3, im not sure what they mean by "features". #4 is asking you to show the steps you took to get: a(x^2 + bx + c)

OpenStudy (adrianna.gongora):

Okay so I would put 1. a(x^2 + bx + c) 2. Polynomial Function 3. 4.The function j(x) can only have a GCF factored out of it. this has form ax^2 + abx + ac factors to a(x^2 + bx + c)

OpenStudy (i<3u):

Yes, I cant figure out what they mean by "features" ;-;

OpenStudy (adrianna.gongora):

okay :) can you help me with one more?

OpenStudy (i<3u):

I need to finish a Biology assignment first. I had only time to help you, then im going to do my work. When im done, ill help you :)

OpenStudy (adrianna.gongora):

Okay :)